Timeline
Mar 16, 2016:
- 8:06 PM Ticket #1006 (16 Mar 19:52 Operation Resume) created by
 - Person in Charge: F. Montenegro. Experiment: EW_drifts+Imaging 1) 08:00 - …
 - 7:53 PM Ticket #1005 (16 Mar 19:38 Figures) created by
 - Figures of rtis and spectras.
 - 7:49 PM Ticket #1004 (16 Mar 19:36 Noise levels) created by
 - 16-03-2016 09:11 EW-DRIFTS A:23.47 B:22.38 C:23.75 D:23.35 12:18 …
 - 7:47 PM Ticket #1003 (16 Mar 19:32 Spark gaps adjustment) created by
 - 16-03-2016 09:15 Spark gaps adjustment 11:38 Spark gaps adjustment 13:40 …
 - 7:16 PM Ticket #1002 (16 Mar 19:00 Transmitter Power Measurements) created by
 - 08:00 TX Power sample (-70 dB coupling) Tx1 +6.4 dBm Tx4 +6.4 dBm Line …
 - 5:55 PM Ticket #1001 (16 Mar 17:35 Imaging Starts) created by
 - 12:55 PM Ticket #834 (25 Jan 23:59 Experiment Plots) closed by
 - invalid
 - 12:25 PM Ticket #1000 (16 Mar 12:00 JASMET30 Manual Configuration) created by
 - 8:05 AM Ticket #999 (16 Mar 08:00 Operation resume) created by
 - Person in Charge: P. Condor. Experiment: EW_drifts+Imaging 1) 20:00 - …
 - 7:53 AM Ticket #998 (16 Mar 07:55 Figures) created by
 - Figures of rtis and spectras.
 - 7:36 AM Ticket #997 (16 Mar 06:08 Noise levels) created by
 - 15-03-2016 21:05 EW-DRIFTS A:22.49 B:21.10 C:22.22 D:21.79 IMAGING …
 - 7:21 AM Ticket #996 (16 Mar 07:00 Transmitter power measurement) created by
 - 15/03/2016 20:00 TX Power sample (-70 dB coupling) Tx1 +6.5 dBm Tx4 +6.5 …
 - 7:12 AM Ticket #995 (16 Mar 01:36 Imaging off) created by
 - Imaging off.
 - 7:10 AM Ticket #994 (16 Mar 06:32 Spark gaps adjustment) created by
 - 15 -03-2016 20:20 Spark gaps adjustment 22:57 Spark gaps adjustment …
 - 7:04 AM Ticket #993 (15 Mar 21:04 Transmitter is On) created by
 - Transmitter is On
 - 7:01 AM Ticket #992 (15 Mar 20:50 Problem with transmitter) created by
 - Problem with transmitter
 - 12:28 AM Ticket #991 (16 Mar 00:10 Figures) created by
 - Figures of rtis and spectras.
 
Mar 15, 2016:
- 8:07 PM Ticket #990 (15 Mar 19:55 Operation Resume) created by
 - Person in Charge: R. Yanque Experiment: EW_drifts+Imaging 1) 08:00 - 20:00 …
 - 8:00 PM Ticket #989 (15 Mar 19:25 Spread F Starts) created by
 - 7:59 PM Ticket #988 (15 Mar 19:00 Noise Levels) created by
 - 13:00 ChA 21.30 ChB 21.97 ChC 22.83 ChD 22.44 16:29 ChA …
 - 7:53 PM Ticket #987 (15 Mar 17:35 Spark Gap Adjustments) created by
 - 08:42 Spark Gap Adjustment 10:42 Spark Gap Adjustment 12:00 Spark Gap …
 - 7:50 PM Ticket #986 (15 Mar 17:52 Experiment stopped) created by
 - The experiment was stopped to repair the Channel B of Imaging. The TR of …
 - 7:48 PM Ticket #985 (15 Mar 17:36 Imaging on) created by
 - Imaging started with Channel B with lower noise 7 dB, less than the other …
 - 7:47 PM Ticket #984 (15 Mar 19:30 Transmitter power measurement) created by
 - 08:00 TX Power sample (-70 dB coupling) Tx1 +6.5 dBm Tx4 +6.5 dBm Line …
 - 7:46 PM Ticket #983 (15 Mar 16:19 Acquisition stopped) created by
 - Acquisition stopped due to PC stopped. It restarted at 16:51
 - 7:45 PM Ticket #982 (15 Mar 14:28 Acquisition stopped) created by
 - The acquisition stopped due to the PC stopped. It was restarted at 14:42
 - 7:43 PM Ticket #981 (15 Mar 12:58 Lower Power) created by
 - Due to failure in Driver. It was fixed at 13:06
 - 7:41 PM Ticket #980 (15 Mar 11:42 Transmitter Failure) created by
 - The transmitter was on at 11:58
 - 9:11 AM Ticket #978 (15 Mar 07:50 SPARK GAPS ADJUSTMENT) closed by
 - invalid
 - 9:03 AM Ticket #975 (23:59 14 Mar 2016 FIgures) closed by
 - invalid
 - 9:03 AM Ticket #974 (23:59 14 Mar 2016 FIgures) closed by
 - invalid
 - 8:17 AM Ticket #979 (15 Mar 07:59 OPERATION RESUME) created by
 - Person in Charge: A. Valdez. Experiment: EW_drifts+Imaging 1) 20:00 - …
 - 8:05 AM Ticket #978 (15 Mar 07:50 SPARK GAPS ADJUSTMENT) created by
 - 14-03-2016 21:00 Spark gaps adjustment 15-03-2016 00:10 Spark gaps …
 - 8:04 AM Ticket #977 (15 Mar 07:50 SPARK GAPS ADJUSTMENT) created by
 - 14-03-2016 21:00 Spark gaps adjustment 15-03-2016 00:10 Spark gaps …
 - 8:01 AM Ticket #976 (14 Mar 23:59 FIgures) created by
 - 8:00 AM Ticket #975 (23:59 14 Mar 2016 FIgures) created by
 - 8:00 AM Ticket #974 (23:59 14 Mar 2016 FIgures) created by
 - 7:13 AM Ticket #973 (15 Mar 07:00 Transmitter Power Measurements) created by
 - 00:00 TX Power sample (-70 dB coupling) Tx1 +6.5 dBm Tx4 +6.6 dBm D3 - 6.8 …
 - 12:18 AM Ticket #972 (15 Mar 00:05 Transmitter Power Measurements) created by
 - 20:00 TX Power sample (-70 dB coupling) Tx1 +6.5 dBm Tx4 +6.6 dBm D3 - 6.9 …
 
Mar 14, 2016:
- 8:14 PM Ticket #971 (14 Mar 19:59 Operation Resume) created by
 - Person in Charge: I. Manay. Experiment: ISR + Imaging 1) 18:30 - 20:00 2) …
 - 8:10 PM Ticket #970 (14 Mar 19:51 Re-start of ew_drift experiment) created by
 - Ew_drift experiment was restarted due to hung up of Ares PC.
 - 7:35 PM Ticket #969 (14 Mar 19:00 Experiments were re-established) created by
 - The B field amplifier of the Imaging experiment was changed due to fall of …
 - 7:26 PM Ticket #968 (14 Mar 18:30 ew_drift and Imaging experiments started) created by
 - System was run with the configuration of february due to interferences …
 - 11:45 AM Ticket #967 (14 Mar 11:30 Antenna VSWR) created by
 - Antenna VSWR WUp 1.07 WDn 1.08 EUp 1.05 EDn 1.07 Hysell 1.11 …
 - 10:13 AM Ticket #966 (14 Mar 10:00 ISCOD Log Created) created by
 - ISCOD MEridian Circle EW Drift - Imaging L. User: Percy Condor Start: ASAP …
 - 10:07 AM Ticket #866 (28 Jan 20:00 Operation Resume) closed by
 - invalid
 - 9:50 AM Ticket #875 (29 Jan 23:59 Plots of Experiment) closed by
 - invalid
 - 9:49 AM Ticket #884 (30 Jan 19:59 Figures) closed by
 - invalid
 - 9:44 AM Ticket #965 (31 Jan 20:00 Plots) created by
 
Feb 18, 2016:
- 2:43 PM Ticket #964 (18 Feb 14:00 Final Report and Kpi) created by
 - The final report
 - 2:43 PM Ticket #963 (18 Feb 14:00 Final Report and Kpi) created by
 - The final report
 - 2:42 PM Ticket #962 (18 Feb 14:00 Final Report and Kpi) created by
 - The final report
 
Feb 16, 2016:
- 12:16 PM Ticket #961 (16 Feb 10:30 Antenna VSWR) created by
 - WUp 1.02 WDn 1.06 NDn 1.05 NUp 1.13 EUp 1.04 EDn 1.03 SDn 1.05 SUp 1.07
 - 12:12 PM Ticket #960 (15 Feb 12:00 Antenna VSWR) created by
 - WUp 1.05 WDn 1.11 NDn 1.09 NUp 1.01 EUp 1.05 EDn 1.01 SDn 1.04 SUp 1.03
 - 10:35 AM Ticket #959 (15 Feb 10:00 Test starts) created by
 - 15 Feb Long Pulse Faraday 3 different long pulses 1ms, 2ms and 4 ms Antena …
 
Note: See TracTimeline
        for information about the timeline view.
    